LASER VELOCITY MODULATION SPECTROSCOPY OF THE ^{3}Δ(3d4s) - X^{3}Φ(3d^{2}) VISIBLE SYSTEM OF TiCl^{+} IN THE VISIBLE RANGE AND CHARACTERIZATION OF THE SPIN-ORBIT STRUCTURE.

C. FOCSA, C. DUFOUR, B. PINCHEMEL, Laboratoire de Dynamique Moleculaire et Photonique, Centre d’Etudes et de Recherches Lasers et Applications, Universite de Lille, 59 655 Villeneuve d’Ascq cedex, France; I. HADJ BACHIR and \underline{T. R. HUET}, Laboratoire de Spectroscopie Hertzienne, Centre d’Etudes et de Recherches Lasers et Applications, Universite de Lille, 59 655 Villeneuve d’Ascq cedex, France..

A single mode cw dye laser along with velocity modulation detection was used to record the absorption spectrum of the the visible system of TiCl^{+} produced in an ac glow discharge between 17 100 and 18 600 cm^{−1} with a gas mixture of He/TiCl_{4}. The translational temperature can be estimated to be about 800 K inducing narrow line shape. The rotational structure of the (0,0) and (1,0) vibrational bands has been observed and fully analysed for the main isotopomer Ti^{35}Cl^{+} as well as the (0,0) and (1,0) bands for Ti^{37} Cl^{+}. A set of molecular parameters has been obtained which includes the spin-orbit parameters thanks to the observation of the forbidden ^{3}Δ_{2} - ^{3}Φ_{2} and ^{3}Δ_{3} - ^{3}Φ_{3} intercombination transitions^{a}. The X^{3}Φ(3d^{2}) state was found to be perturbed. Additional spectra have been recorded with a better sensitivity in the same spectral range. In this communication we will report the observation of new bands. Among them, hot bands have already been identified, involving the v′′ = 1 levels. The detailed analysis of the spectrum in this range is currently performed in order to determine a set of unperturbed molecular parameters for the ground state. Efforts are also in progress for producing other Ti - containing ions by introduction of a reactive gas into the discharge, in relation with the astrophysical interest of Ti containing compounds.

aC. Focsa, C. Dufour, B. Pinchemel, I. Hadj Bachir, and T.R. Huet J. Chem. Phys. (submitted for publication)